# Complete non-ambiguous trees and associated permutations: new enumerative results

### v2

@@ 4.1. Linking permutations with one and two CNATs
In this part, we establish a bijection between marked permutations with a single CNAT, and permutations with $2$ CNATs (Theorem 4.3). We begin by describing the set $B(n,1)$ of permutations with a single CNAT. In [11], the authors gave a characterisation of this set in terms of so-called L-subsets. We first reformulate their characterisation in terms of quadrants. For an (irreducible) $n$-permutation $\pi$, we let $\Pi=\{(i,\pi_{i});i\in[n]\}$ be the points in the plot of $\pi$. Given an index $k\in\{2,\cdots,n\}$, we partition $\Pi$ into four quadrants, called $k$-quadrants when the index $k$ needs to be explicit, as follows.
@@
This partition is illustrated on Figure 7 below, with $\pi=561243$ and $k=4$.
@@
Note that since $\pi$ is a permutation, the first $k-1$ columns or rows of its plot must have $k-1$ dots in total. Therefore we have
@@
so that $\left|\Pi_{<k,\geq k}\right|=\left|\Pi_{\geq k,<k}\right|$. Moreover, this number must be strictly positive, since otherwise $\pi$ would induce a permutation on $[k-1]$, and therefore be reducible. In words, given an (irreducible) $n$-permutation $\pi$, and an index $k\in\{2,\cdots,n\}$, the lower-left and upper-right quadrants are non-empty, and have the same number of dots.
@@
(3) $\forall k\in\{2,\cdots,n\},\ \left|\Pi_{<k,\geq k}\right|=\left|\Pi_{\geq k,<k}\right|=1.$
@@
Let $\pi$ be an $n$-permutation. For $j\in[n]$, we define the $j$-th L-subset $L_{j}$ of a permutation $\pi$ by $L_{j}:=\Pi_{<j+1,<j+1}\setminus\Pi_{<j,<j}$. Therorem 3.11 in [11] states that a permutation has a unique CNAT if, and only if, $\pi$ has no fixed point, $|L_{1}|=0$, $|L_{n}|=2$, and $|L_{j}|=1$ for all $j\in\{2,\cdots,n-1\}$. We show that this is equivalent to $\pi$ having no fixed point and satisfying the quadrant condition.
@@
Suppose first that $\pi$ satisfies the above conditions on its L-subsets, and fix some $k\in\{2,\cdots,n\}$. Then we have $\left|\Pi_{<k,<k}\right|=\sum\limits_{j=1}^{k-1}|L_{j}|=k-2$. By Equation (2) this implies that $\left|\Pi_{<k,\geq k}\right|=\left|\Pi_{\geq k,<k}\right|=k-1-\left|\Pi_{<k,<k}\right|=1$. Thus $\pi$ satisfies the quadrant condition as desired.
@@
Conversely, suppose that $\pi$ satisfies the quadrant condition, and has no fixed point. Owing to the fact that $\pi$ is irreducible, we have $|L_{1}|=0$ (we cannot have $\pi_{1}=1$). Moreover there is always exactly one dot in the rightmost column $n$ and exactly one dot in the bottom-most row $n$, which both belong to the L-subset $L_{n}$. Again, $\pi$ is irreducible, so these dots must be distinct (we cannot have $\pi_{n}=n$, leading to the equality $|L_{n}|=2$. It therefore remains to show that
(4) $\forall j\in\{2,\cdots,n-1\},\,|L_{j}|=1.$
@@
Seeking contradiction, suppose that this is not the case. Because $\pi$ is a permutation, there are $n$ dots in total in $\Pi$, and therefore $(n-2)$ dots in $\bigcup\limits_{j=2}^{n-1}L_{j}$. Thus if the L-condition (4) is not satisfied, there must exist $j\in\{2,\cdots,n-1\}$ such that the $j$-th L-subset $L_{j}$ is empty. In that case, the dot in the column $j$ will be placed below row $j$, i.e. in column $j$ of the lower-left quadrant $\Pi_{<j+1,\geq j+1}$. Now since $L_{j}$ is empty, each dot located in the first $(j-1)$ columns should be placed either in the upper-left quadrant $\Pi_{<j,<j}$ or the lower-left quadrant $\Pi_{<j+1,\geq j+1}$ but before the column $j$. Because $\pi$ is irreducible, there must be at most $(j-2)$ dots in the upper-left quadrant $\Pi_{<j,<j}$, which means that there must be at least one dot placed in the first $(j-1)$ columns of the lower-left quadrant $\Pi_{<j+1,\geq j+1}$, say in some column $j^{\prime}<j$. Hence, in the lower-left quadrant $\Pi_{<j+1,\geq j+1}$ there are at least two dots: one in column $j^{\prime}$ and one in column $j$. This contradicts the fact that $\pi$ satisfies the quadrant condition. Therefore, the L-condition (4) is satisfied, and $\pi$ has a unique CNAT by [11, Theorem 3.11]. ∎
@@
For $n\geq 1$, and an index $j\in\{2,\cdots,n\}$, we now define the insertion operation $\mathrm{Insert}_{[}j]:S_{n}\rightarrow S_{n+1}$ as follows. Given a $n$-permutation $\pi$, we construct $\pi^{\prime}:=\mathrm{Insert}_{[}j](\pi)$ by inserting a fixed point $j$ into the $n$-permutation $\pi$, and increasing the labels of all letters $k\geq j$ in the original permutation $\pi$ by $1$. For example, if $\pi=521634$ and $j=3$, then $\pi^{\prime}=62\textbf{3}1745$ (fixed point in bold). It is straightforward to check that $\mathrm{Insert}_{[}j](\pi)$ is irreducible if, and only if, $\pi$ is irreducible.
@@
Let $n\geq 2$. Define a map $\Phi:\{2,\cdots,n\}\times S_{n}\rightarrow S_{n+1}$ by $\Phi(j,\pi):=\mathrm{Insert}_{[}j](\pi)$ for all $j\in\{2,\cdots,n\}$ and $\pi\in S_{n}$. Then $\Phi$ is a bijection from $\{2,\cdots,n\}\times B(n,1)$ to $B(n+1,2)$.
@@
The insertion index $j\in\{2,\cdots,n\}$ can be thought of as marking the “space” between $\pi_{j-1}$ and $\pi_{j}$ in $\pi$, with the insertion operation corresponding to inserting a fixed point in the marked space. The product set $\{2,\cdots,n\}\times B(n,1)$ can then be thought of as permutations with a single CNAT, marked in such a space. Hence the bijection of Theorem 4.3 can be interpreted as a bijection between marked $n$-permutations with a single CNAT, and $(n+1)$-permutations with two CNATs.
From [11, Corollary 3.12], we have $b(n,1)=2^{n-2}$ for any $n\geq 2$. This formula is also implicit from the combination of Theorem 2.12 and [1, Theorem 1]. Indeed, in that work the authors show that the number of $n$-permutations whose graphs are trees is $2^{n-2}$, and trees are the only simple graphs satisfying $t_{G}=1$ (Lemma 2.4). Together with Theorem 4.3, this enumeration formula immediately implies the following, which answers in the affirmative the conjecture from [11] that $\big{(}b(n,2)\big{)}_{n\geq 2}$ is given by Sequence A001787 in [17].
@@
Let $n\geq 2$, and $j\in\{2,\cdots,n\}$. For any $k\geq 1$, if $\pi$ is an $n$-permutation with $k$ CNATs, then $\mathrm{Insert}_{[}j](\pi)$ is an $(n+1)$-permutation with at least $2k$ CNATs. Moreover, if $\pi$ has a single CNAT (i.e. $k=1$), then $\mathrm{Insert}_{[}j](\pi)$ has exactly $2$ CNATs.
@@
The above construction is illustrated in Figure 8 below. Here we start with a CNAT $T$ and associated permutation $\pi$ (Figure 7(a)). We then show in Figures 7(c) and 7(d) possible choices for the two CNATs $T_{1}$ and $T_{2}$ with permutation $\pi^{\prime}=\mathrm{Insert}_{[}4](\pi)$. Note that here in each case we actually had two choices for the new edge in the CNATs $T_{1}$ and $T_{2}$, since there were two edges of $T$ crossing row $4$ to the left of the new leaf dot (in columns $1$ and $2$), and two edges crossing column $4$ above the new leaf dot (in rows $2$ and $3$).
111122223333444455556666111122223333444455556666
@@
Let $\pi$ be an $n$-permutation, and $k\in\{2,\cdots,n\}$ an index. We show that if the lower-left and upper-right $k$-quadrants have at least $2$ points, then we can construct at least $3$ CNATs with associated permutation $\pi$. The construction is similar to the one in the proof of Lemma 4.6, so we allow ourselves to be a little briefer here.
Let $d_{1},d_{1}^{\prime}$, resp. $d_{2},d_{2}^{\prime}$, be two dots in the lower-left, resp. upper-right, quadrant. We assume that the column of $d_{1}$ is to the left of that of $d_{1}^{\prime}$, and the row of $d_{2}$ above that of $d_{2}^{\prime}$. Let $\pi^{\prime}$ be the permutation obtained from $\pi$ by removing the rows and columns containing $d_{1}^{\prime}$ and $d_{2}^{\prime}$ from its plot. Let $T$ be a CNAT with permutation $\pi^{\prime}$ [^2] See Remark 4.8 for why the assumption that $T$ exists can be justified.. As in the proof of Lemma 4.6, we let $T^{\prime}$ be the “partial” CNAT obtained from $T$ by re-inserting the leaf dots $d_{1}^{\prime}$ and $d_{2}^{\prime}$.
Let $c=(i^{\prime},j^{\prime})$ denote the cell in the same column $i^{\prime}$ as $d_{1}^{\prime}$ and row $j^{\prime}$ as $d_{2}^{\prime}$. By construction this cell is in the upper-left $k$-quadrant of $\pi$. Since $d_{1}$ is in some column $i<i^{\prime}$ (to the left of $d_{1}^{\prime}$), the path from $d_{1}$ to the root in the partial CNAT $T^{\prime}$ must cross the row $j^{\prime}$ in some cell $c_{\mathrm{left}}$ to the left of $c$. Similarly, since the row $j$ of $d_{2}$ is above the row $j^{\prime}$ of $d_{2}^{\prime}$, the path from $d_{2}$ to the root must cross the column $i^{\prime}$ in some cell $c_{\mathrm{up}}$ above $c$. Then putting internal dots in any two of the three cells $c,c_{\mathrm{left}},c_{\mathrm{up}}$ will yield a CNAT with permutation $\pi$. Since there are three such possibilities, this implies that $\pi$ has at least three CNATs, which completes the proof. These constructions are illustrated on Figure 9 below. ∎
@@
(a) The permutation $\pi=461523$. The lower-left and upper-right $4$-quadrants have $2$ dots.
111122223333444455556666111122223333444455556666
@@
Figure 9. Illustrating how to construct three CNATs with an associated permutation $\pi=561423$ which does not satisfy the quadrant condition for $k=4$.
Remark 4.8.
There is a slight imprecision in the proof of Lemma 4.7, when we assume that the permutation $\pi^{\prime}$ obtained by deleting $d_{1}^{\prime}$ and $d_{2}^{\prime}$ from $\pi$ has a CNAT, which is equivalent to $\pi^{\prime}$ being irreducible, or equivalently to the graph $G^{\prime}:=G_{\pi^{\prime}}$ being connected. This may in fact not always be the case (see example below). Denote $(i,\pi_{i})$ the cell of the dot $d_{1}^{\prime}$, and suppose that deleting $\pi_{i}$ from $i$ (and re-labelling other elements) yields a reducible permutation $\pi^{1}\in S_{n-1}$. This means that $\pi^{1}$ has a breakpoint $j$ for some $11\leq j<n-$ 1. Moreover, the permutation $\pi^{2}:=(\pi_{i}-j)(\pi_{j+1}-j)(\pi_{j+2}-j)\cdots(\pi_{n}-j)$ must be irreducible (a breakpoint of $\pi^{2}$ would also be a breakpoint of $\pi$), and as such has a CNAT, say $T^{\prime\prime}$. Now in the CNAT construction, when we delete $d_{1}^{\prime}$, we also delete all dots in columns $j+1$ to $n$. Then, when re-inserting $d_{1}^{\prime}$ into the partial CNAT $T^{\prime}$, we re-insert those columns by “gluing” the CNAT $T^{\prime\prime}$ into the (full) CNATs. The process is the same when removing $d_{2}^{\prime}$, and the rest of the proof is unchanged.
We illustrate this case on Figure 10, with the permutation $\pi=4612375$, and $k=4$. The green dots to delete are $d_{1}^{\prime}$ in cell $(2,6)$ and $d_{2}^{\prime}$ in cell $(5,3)$. However, in this case, deleting the dot $d_{1}^{\prime}$ yields $\pi^{\prime}=(4123)(65)$ which has a breakpoint at $j=4$, so does not have a CNAT. We therefore will also need to delete the dots in columns $6$ and $7$ (in pink). We then construct a partial CNAT $T^{\prime}$ with associated “partial permutation” $412$, resulting from deleting these alongside $d_{1}^{\prime}$ and $d_{2}^{\prime}$, and a partial CNAT $T^{\prime\prime}$ with associated “partial permutation” $675$ (in pink on Figure 9(b)). We then simply “glue” $T^{\prime\prime}$ into the full CNATs that we construct when re-inserting $d_{1}^{\prime}$ and $d_{2}^{\prime}$ as in the general case.
@@
Figure 10. Illustrating how to construct three CNATs with an associated permutation $\pi=461235$.
@@
Let $n\geq 2$, and $\pi\in B(n,2)$ an $n$-permutation with exactly $2$ CNATs. Then $\pi$ has a unique fixed point $j\in\{2,\cdots,n-1\}$.
@@
Much like Proposition 4.2 for $B(n,1)$, Lemmas 4.7 and 4.9 in fact give a characterisation of the sets $B(n,2)$, as follows.
@@
Let $\pi$ be a permutation. Then $\pi$ has exactly $2$ CNATs if, and only if, $\pi$ satisfies the quadrant condition and has a unique fixed point $j\in\{2,\cdots,|\pi|-1\}$.
@@
If $\pi$ has exactly $2$ CNATs, then it satisfies the quadrant condition by Lemma 4.7 and has a unique fixed point $j\in\{2,\cdots,|\pi|-1\}$ by Lemma 4.9. Conversely, if $\pi$ satisfies the quadrant condition and has a unique fixed point $j$, then deleting $j$ from $\pi$ yields a permutation $\pi^{\prime}$ with no fixed point which also satisfies the quadrant condition. Thus $\pi^{\prime}$ has a unique CNAT by Proposition 4.2. Since $\pi=\mathrm{Insert}_{[}j](\pi^{\prime})$ by construction, we deduce that $\pi$ has exactly $2$ CNATs by Lemma 4.6. ∎
Theorem 4.3 now follows straightforwardly from Propositions 4.2 and 4.10. Indeed, we have shown that permutations with a single CNAT are those which satisfy the quadrant condition and have no fixed point, while permutations with two CNATs are those that satisfy the quadrant condition and have a single fixed point. Since the insertion or deletion of a fixed point does not affect whether a permutation satisfies the quadrant condition or not, it is immediate that the map $\Phi$ defined by $\Phi(j,\pi):=\mathrm{Insert}_{[}j](\pi)$ is indeed a bijection from $\{2,\cdots,n\}\times B(n,1)$ to $B(n+1,2)$.

### v4

@@ 4.1. Linking permutations with one and two CNATs
In this part, we establish a bijection between marked permutations with a single CNAT, and permutations with $2$ CNATs (Theorem 4.3). We begin by describing the set $B(n,1)$ of permutations with a single CNAT. In [11], the authors gave a characterisation of this set in terms of so-called L-subsets. We first reformulate their characterisation in terms of quadrants. For an (irreducible) $n$-permutation $\pi$, we let $\Pi=\{(i,\pi_{i});i\in[n]\}$ be the points in the plot of $\pi$. Given an index $k\in\{2,\ldots,n\}$, we partition $\Pi$ into four quadrants, called $k$-quadrants when the index $k$ needs to be explicit, as follows.
@@
This partition is illustrated in Figure 7 below, with $\pi=561243$ and $k=4$.
@@
Note that since $\pi$ is a permutation, the first k1k-1 columns or rows of its plot must have k1k-1 dots in total. Therefore we have
@@
so that $\left|\Pi_{<k,\geq k}\right|=\left|\Pi_{\geq k,<k}\right|$. Moreover, this number must be strictly positive, since otherwise $\pi$ would induce a permutation on [k1][k-1], and therefore be reducible. In words, given an (irreducible) $n$-permutation $\pi$, and an index $k\in\{2,\ldots,n\}$, the lower-left and upper-right quadrants are non-empty, and have the same number of dots.
@@
(3) $\forall k\in\{2,\ldots,n\},\ \left|\Pi_{<k,\geq k}\right|=\left|\Pi_{\geq k,<k}\right|=1.$
@@
Let $\pi$ be an $n$-permutation. For $j\in[n]$, we define the $j$-th L-subset $L_{j}$ of a permutation $\pi$ by $L_{j}:=\Pi_{<j+1,<j+1}\setminus\Pi_{<j,<j}$. Therorem 3.11 in [11] states that a permutation has a unique CNAT if, and only if, $\pi$ has no fixed point, $|L_{1}|=0$, $|L_{n}|=2$, and $|L_{j}|=1$ for all $j\in\{2,\ldots,n-1\}$. We show that this is equivalent to $\pi$ having no fixed point and satisfying the quadrant condition.
@@
Suppose first that $\pi$ satisfies the above conditions on its L-subsets, and fix some $k\in\{2,\ldots,n\}$. Then we have $\left|\Pi_{<k,<k}\right|=\sum\limits_{j=1}^{k-1}|L_{j}|=k-2$. By Equation (2) this implies that $\left|\Pi_{<k,\geq k}\right|=\left|\Pi_{\geq k,<k}\right|=k-1-\left|\Pi_{<k,<k}\right|=1$. Thus $\pi$ satisfies the quadrant condition as desired.
@@
Conversely, suppose that $\pi$ satisfies the quadrant condition, and has no fixed point. Owing to the fact that $\pi$ is irreducible, we have $|L_{1}|=0$ (we cannot have $\pi_{1}=1$). Moreover there is always exactly one dot in the rightmost column $n$ and exactly one dot in the bottom-most row $n$, which both belong to the L-subset $L_{n}$. Again, $\pi$ is irreducible, so these dots must be distinct (we cannot have $\pi_{n}=n$), leading to the equality $|L_{n}|=2$. It therefore remains to show that
(4) $\forall j\in\{2,\ldots,n-1\},\,|L_{j}|=1.$
@@
Seeking contradiction, suppose that this is not the case. Because $\pi$ is a permutation, there are $n$ dots in total in $\Pi$, and therefore (n2)(n-2) dots in $\bigcup\limits_{j=2}^{n-1}L_{j}$. Thus if the L-condition (4) is not satisfied, there must exist $j\in\{2,\ldots,n-1\}$ such that the $j$-th L-subset $L_{j}$ is empty. In that case, the dot in the column $j$ will be placed below row $j$, i.e. in column $j$ of the lower-left quadrant $\Pi_{<j+1,\geq j+1}$. Now since $L_{j}$ is empty, each dot located in the first (j1)(j-1) columns should be placed either in the upper-left quadrant $\Pi_{<j,<j}$ or the lower-left quadrant $\Pi_{<j+1,\geq j+1}$ but before the column $j$. Because $\pi$ is irreducible, there must be at most (j2)(j-2) dots in the upper-left quadrant $\Pi_{<j,<j}$, which means that there must be at least one dot placed in the first (j1)(j-1) columns of the lower-left quadrant $\Pi_{<j+1,\geq j+1}$, say in some column $j^{\prime}<j$. Hence, in the lower-left quadrant $\Pi_{<j+1,\geq j+1}$ there are at least two dots: one in column $j^{\prime}$ and one in column $j$. This contradicts the fact that $\pi$ satisfies the quadrant condition. Therefore, the L-condition (4) is satisfied, and $\pi$ has a unique CNAT by [11, Theorem 3.11]. ∎
@@
For $n\geq 1$, and an index $j\in\{2,\ldots,n\}$, we now define the insertion operation $\mathrm{Insert}_{[}j]:S_{n}\rightarrow S_{n+1}$ as follows. Given a $n$-permutation $\pi$, we construct $\pi^{\prime}:=\mathrm{Insert}_{[}j](\pi)$ by inserting a fixed point $j$ into the $n$-permutation $\pi$, and increasing the labels of all letters $k\geq j$ in the original permutation $\pi$ by $1$. For example, if $\pi=521634$ and $j=3$, then $\pi^{\prime}=62\textbf{3}1745$ (fixed point in bold). It is straightforward to check that $\mathrm{Insert}_{[}j](\pi)$ is irreducible if, and only if, $\pi$ is irreducible.
@@
Let $n\geq 2$. Define a map $\Phi:\{2,\ldots,n\}\times S_{n}\rightarrow S_{n+1}$ by $\Phi(j,\pi):=\mathrm{Insert}_{[}j](\pi)$ for all $j\in\{2,\ldots,n\}$ and $\pi\in S_{n}$. Then $\Phi$ is a bijection from $\{2,\ldots,n\}\times B(n,1)$ to $B(n+1,2)$.
@@
The insertion index $j\in\{2,\ldots,n\}$ can be thought of as marking the “space” between $\pi_{j-1}$ and $\pi_{j}$ in $\pi$, with the insertion operation corresponding to inserting a fixed point in the marked space. The product set $\{2,\ldots,n\}\times B(n,1)$ can then be thought of as permutations with a single CNAT, marked in such a space. Hence the bijection of Theorem 4.3 can be interpreted as a bijection between marked $n$-permutations with a single CNAT, and $(n+1)$-permutations with two CNATs.
From [11, Corollary 3.12], we have $b(n,1)=2^{n-2}$ for any $n\geq 2$. This formula is also implicit from the combination of Theorem 2.13 and [1, Theorem 1]. Indeed, in that work the authors show that the number of $n$-permutations whose graphs are trees is $2^{n-2}$, and trees are the only simple graphs satisfying $a_{G}=1$ (Lemma 2.4). Together with Theorem 4.3, this enumeration formula immediately implies the following, which answers in the affirmative the conjecture from [11] that $\big(b(n,2)\big)_{n\geq 2}$ is given by Sequence A001787 in [16].
@@
Let $n\geq 2$, and $j\in\{2,\ldots,n\}$. For any $k\geq 1$, if $\pi$ is an $n$-permutation with $k$ CNATs, then $\mathrm{Insert}_{[}j](\pi)$ is an $(n+1)$-permutation with at least $2k$ CNATs. Moreover, if $\pi$ has a single CNAT (i.e. $k=1$), then $\mathrm{Insert}_{[}j](\pi)$ has exactly $2$ CNATs.
@@
The above construction is illustrated in Figure 8 below. Here we start with a CNAT $T$ and associated permutation $\pi$ (Figure 8(a)). We then show in Figures 8(c) and 8(d) possible choices for the two CNATs $T_{1}$ and $T_{2}$ with permutation $\pi^{\prime}=\mathrm{Insert}_{[}4](\pi)$. Note that here in each case we actually had two choices for the new edge in the CNATs $T_{1}$ and $T_{2}$, since there were two edges of $T$ crossing row $4$ to the left of the new leaf dot (in columns $1$ and $2$), and two edges crossing column $4$ above the new leaf dot (in rows $2$ and $3$).
$112233445566$
@@
Let $\pi$ be an $n$-permutation, and $k\in\{2,\ldots,n\}$ an index. We show that if the lower-left and upper-right $k$-quadrants have at least $2$ points, then we can construct at least $3$ CNATs with associated permutation $\pi$. The construction is similar to the one in the proof of Lemma 4.6, so we allow ourselves to be a little briefer here.
Let $d_{1},d_{1}^{\prime}$, resp. $d_{2},d_{2}^{\prime}$, be two dots in the lower-left, resp. upper-right, quadrant. We assume that the column of $d_{1}$ is to the left of that of $d_{1}^{\prime}$, and the row of $d_{2}$ above that of $d_{2}^{\prime}$. Let $\pi^{\prime}$ be the permutation obtained from $\pi$ by removing the rows and columns containing $d_{1}^{\prime}$ and $d_{2}^{\prime}$ from its plot. Let $T$ be a CNAT with permutation $\pi^{\prime}$ [^2] See the proof of Lemma 4.7 for why the assumption that $T$ exists can be justified.. As in the proof of Lemma 4.6, we let $T^{\prime}$ be the “partial” CNAT obtained from $T$ by re-inserting the leaf dots $d_{1}^{\prime}$ and $d_{2}^{\prime}$.
Let $c=(i^{\prime},j^{\prime})$ denote the cell in the same column $i^{\prime}$ as $d_{1}^{\prime}$ and row $j^{\prime}$ as $d_{2}^{\prime}$. By construction this cell is in the upper-left $k$-quadrant of $\pi$. Since $d_{1}$ is in some column $i<i^{\prime}$ (to the left of $d_{1}^{\prime}$), the path from $d_{1}$ to the root in the partial CNAT $T^{\prime}$ must cross the row $j^{\prime}$ in some cell $c_{\mathrm{left}}$ to the left of $c$. Similarly, since the row $j$ of $d_{2}$ is above the row $j^{\prime}$ of $d_{2}^{\prime}$, the path from $d_{2}$ to the root must cross the column $i^{\prime}$ in some cell $c_{\mathrm{up}}$ above $c$. Then putting internal dots in any two of the three cells $c,c_{\mathrm{left}},c_{\mathrm{up}}$ will yield a CNAT with permutation $\pi$. Since there are three such possibilities, this implies that $\pi$ has at least three CNATs, which completes the proof. These constructions are illustrated in Figure 9 below.
@@
(a) The permutation $\pi=461523$. The lower-left and upper-right $4$-quadrants have 22 dots.
$112233445566$
@@
Figure 9. Illustrating how to construct three CNATs with an associated permutation $\pi=461523$ which does not satisfy the quadrant condition for $k=4$.
However, there is a slight imprecision in the above argument, when we assume that the permutation $\pi^{\prime}$ obtained by deleting $d_{1}^{\prime}$ and $d_{2}^{\prime}$ from $\pi$ has a CNAT, which is equivalent to $\pi^{\prime}$ being irreducible, or equivalently to the graph $G^{\prime}:=G_{\pi^{\prime}}$ being connected. This may in fact not always be the case (see example below). Denote $(i,\pi_{i})$ the cell of the dot $d_{1}^{\prime}$, and suppose that deleting $\pi_{i}$ from $\pi$ (and re-labelling other elements) yields a reducible permutation $\pi^{1}\in S_{n-1}$. This means that $\pi^{1}$ has a breakpoint $j$ for some $1\leq j<n-1$. Moreover, the permutation $\pi^{2}:=(\pi_{i}-j)(\pi_{j+1}-j)(\pi_{j+2}-j)\cdots(\pi_{n}-j)$ must be irreducible (a breakpoint of $\pi^{2}$ would also be a breakpoint of $\pi$), and as such has a CNAT, say $T^{\prime\prime}$. Now in the CNAT construction, when we delete $d_{1}^{\prime}$, we also delete all dots in columns $j+1$ to $n$. Then, when re-inserting $d_{1}^{\prime}$ into the partial CNAT $T^{\prime}$, we re-insert those columns by “gluing” the CNAT $T^{\prime\prime}$ into the (full) CNATs. The process is the same when removing $d_{2}^{\prime}$, and the rest of the proof is unchanged.
We illustrate this case in Figure 10, with the permutation $\pi=4612375$, and $k=4$. The green dots to delete are $d_{1}^{\prime}$ in cell $(2,6)$ and $d_{2}^{\prime}$ in cell $(5,3)$. However, in this case, deleting the dot $d_{1}^{\prime}$ yields $\pi^{\prime}=(4123)(65)$ which has a breakpoint at $j=4$, so does not have a CNAT. We therefore will also need to delete the dots in columns $6$ and $7$ (in pink). We then construct a partial CNAT $T^{\prime}$ with associated “partial permutation” $412$, resulting from deleting these alongside $d_{1}^{\prime}$ and $d_{2}^{\prime}$, and a partial CNAT $T^{\prime\prime}$ with associated “partial permutation” $675$ (in pink in Figure 10(b)). We then simply “glue” $T^{\prime\prime}$ into the full CNATs that we construct when re-inserting $d_{1}^{\prime}$ and $d_{2}^{\prime}$ as in the general case.
@@
Figure 10. Illustrating how to construct three CNATs with an associated permutation $\pi=4612375$.
@@
Let $n\geq 2$, and $\pi\in B(n,2)$ an $n$-permutation with exactly $2$ CNATs. Then $\pi$ has a unique fixed point $j\in\{2,\ldots,n-1\}$.
@@
Much like Proposition 4.2 for $B(n,1)$, Lemmas 4.7 and 4.8 in fact give a characterisation of the set $B(n,2)$, as follows.
@@
Let $\pi$ be a permutation. Then $\pi$ has exactly $2$ CNATs if, and only if, $\pi$ satisfies the quadrant condition and has a unique fixed point $j\in\{2,\ldots,n-1\}$.
@@
If $\pi$ has exactly $2$ CNATs, then it satisfies the quadrant condition by Lemma 4.7 and has a unique fixed point $j\in\{2,\ldots,n-1\}$ by Lemma 4.8. Conversely, if $\pi$ satisfies the quadrant condition and has a unique fixed point $j$, then deleting $j$ from $\pi$ yields a permutation $\pi^{\prime}$ with no fixed point which also satisfies the quadrant condition. Thus $\pi^{\prime}$ has a unique CNAT by Proposition 4.2. Since $\pi=\mathrm{Insert}_{[}j](\pi^{\prime})$ by construction, we deduce that $\pi$ has exactly $2$ CNATs by Lemma 4.6. ∎
Theorem 4.3 now follows straightforwardly from Propositions 4.2 and 4.9. Indeed, we have shown that permutations with a single CNAT are those which satisfy the quadrant condition and have no fixed point, while permutations with two CNATs are those that satisfy the quadrant condition and have a single fixed point. Since the insertion or deletion of a fixed point does not affect whether a permutation satisfies the quadrant condition or not, it is immediate that the map $\Phi$ defined by $\Phi(j,\pi):=\mathrm{Insert}_{[}j](\pi)$ is indeed a bijection from $\{2,\ldots,n\}\times B(n,1)$ to $B(n+1,2)$.